please reply soon

Grade:10

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear deepak sharma,

Chare Q= C*V

Charge,   Q = C × V   where: Q = charge in coulombs (C)
C = capacitance in farads (F)
V = voltage in volts (V)

 

due to this charge there is enrgy also stored. that energy is

Energy,   E = ½QV = ½CV²

 


Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay